OpenStudy (anonymous):

Is cosx=0.61 find sin (x-pie over 2)

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

mmm pie \(\pi\)

OpenStudy (misty1212):

\[\sin(x-\frac{\pi}{2})=-\cos(x)\]

OpenStudy (misty1212):

so put a minus sign in front of \(.61\)

OpenStudy (anonymous):

No , it's if cos X is equal to 0.61 t6th en find sin and it gives me the X minus pie over 2

OpenStudy (anonymous):

*then

OpenStudy (misty1212):

yeah i got the question

OpenStudy (misty1212):

if \[\cos(x)=.61\] then what is \[\sin(x-\frac{\pi}{2})\]

OpenStudy (misty1212):

it is always true, no matter what \(x\) is, that \[\sin(x-\frac{\pi}{2})=-\cos(x)\]

OpenStudy (misty1212):

so if \(\cos(x)=.61\) then \(\sin(x-\frac{\pi}{2})=-\cos(x)=-.61\)